# Airfoil Diffusion: Denoising Diffusion Model For Conditional Airfoil Generation

### v2

@@ Abstract
The design of aerodynamic shapes, such as airfoils, has traditionally required significant computational resources and relied on predefined design parameters, which limit the potential for novel shape synthesis. In this work, we introduce a data-driven methodology for airfoil generation using a diffusion model. Trained on a dataset of preexisting airfoils, our model can generate an arbitrary number of new airfoils from random vectors, which can be conditioned on specific aerodynamic performance metrics such as lift ($C_{l}$) or drag ($C_{d}$). Our results demonstrate that the diffusion model effectively produces airfoil shapes with realistic aerodynamic properties, offering substantial improvements in efficiency, flexibility, and the potential for discovering innovative airfoil designs. This approach significantly expands the design space, facilitating the synthesis of high-performance aerodynamic shapes that transcend the limitations of traditional methods.
@@ 1 Introduction
The design and optimization of aerodynamic shapes, particularly airfoils, are fundamental challenges in aerospace engineering. Traditionally, the process of designing airfoils has relied heavily on computationally intensive simulations to explore the vast design space [1, 2]. This approach typically involves defining specific design parameters that can limit the ability to synthesize truly novel shapes [3]. As the demand for higher performance and more efficient aerodynamic structures increases, there is a growing need for advanced methodologies that can streamline the design process and facilitate the generation of innovative airfoil geometries [4].
Recent advances in machine learning have opened new avenues for the design and optimization of complex systems [5]. Data-driven approaches, such as generative models, have shown promise in various fields by learning from existing data to generate new high-quality samples[6, 7]. In the context of airfoil design, these models can potentially revolutionize the way we approach shape optimization by learning the underlying distribution of airfoil geometries and their aerodynamic properties [3, 8, 9].
@@
The diffusion model, once trained, allows for the generation of new airfoil shapes from random vectors. This capability enables the synthesis of new airfoils that can be selected on the basis of desired Cl or Cd values. Our approach offers a data-driven framework that not only enhances the efficiency of the airfoil design process, but also has the potential to discover novel shapes with superior aerodynamic performance.
@@ 2.1 Geometry Parameterization
Traditional airfoil design has predominantly relied on computational simulations and fixed design parameters, which can restrict the exploration of innovative aerodynamic shapes [10]. These conventional methods, though effective in specific scenarios, often limit the ability to fully explore the expansive design space necessary for optimizing aerodynamic structures [4]. This limitation arises primarily from the need for predefining the design space using specific shape parameters, such as camber and thickness, which can constrain the range of possible geometries that can be considered during optimization.
To address these limitations, there has been a shift towards advanced machine learning techniques that allow for more flexible and efficient exploration of the design space. These approaches, including techniques like Bézier and B-spline parameterizations, offer greater control over airfoil geometry by representing shapes with a set of control points that can be adjusted to generate smooth and continuous surfaces. Such methods have been integrated into optimization frameworks to enhance the precision of geometric design while maintaining smoothness and manufacturability [11, 12]. However, while these traditional geometric parameterizations provide a level of flexibility, they still require significant manual intervention and are often limited in handling highly complex shapes, especially in high-dimensional design spaces.
@@
Moreover, Diffusion Models, as explored in recent studies, represent a significant advancement in airfoil geometry parameterization. These models operate within a learned latent space, allowing for the generation of airfoil shapes with fewer training samples and maintaining high-quality outputs [14]. The ability of Diffusion Models to integrate specific design constraints directly into the generative process provides a powerful tool for optimizing aerodynamic shapes in a data-efficient manner.
As the field continues to evolve, hybrid approaches that combine traditional geometric parameterization with advanced machine learning techniques are emerging. These methods aim to leverage the strengths of both worlds—preserving the smoothness and manufacturability of traditional parameterizations while exploiting the flexibility and data-driven insights of deep learning models. The ongoing research in this area focuses on refining these techniques to accommodate increasingly complex aerodynamic and geometric constraints, ultimately pushing the boundaries of what is possible in airfoil design.
@@ 2.2 Deep Learning Approaches
Generative models, including Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs), have shown considerable promise in optimizing airfoil design by overcoming the limitations of traditional design methods. For instance, the Airfoil GAN model [3] encodes existing airfoil shapes into latent vectors, enabling the generation of novel airfoils with specific aerodynamic properties, without relying on predefined design parameters. This approach not only facilitates the exploration of high-dimensional design spaces but also integrates a genetic algorithm for optimizing aerodynamic performance, allowing the synthesized airfoils to evolve towards desired properties. Moreover, the model’s ability to generate smooth and realistic airfoils directly from the latent space, coupled with its interpretability through feature clustering, sets it apart from traditional techniques. In comparison, the use of free-form deformation generalized adversarial networks (FFD-GAN) [8] has advanced the parameterization of three-dimensional aerodynamic shapes, yet it often requires predefined design formulations, limiting its flexibility in exploring novel geometries. The Airfoil GAN’s potential for application to other aerodynamic and engineering domains further underscores its significance in the field of design optimization.
Despite their potential, GAN-based methods often face challenges related to training stability and the need for large datasets. Addressing these issues, DiffAirfoil [15] introduces a novel approach that leverages a diffusion model within a learned latent space, specifically tailored for aerodynamic shape optimization. This model significantly enhances data efficiency, requiring substantially fewer training geometries compared to GANs, while still generating high-quality and diverse airfoil samples. The diffusion process not only ensures stable training and robust sampling, even under data-scarce conditions, but also mitigates the common pitfalls associated with adversarial training, such as mode collapse.
@@ 3.2.2 NeuralFoil Overview
NeuralFoil leverages a deep learning model trained on a comprehensive dataset of airfoil shapes and their corresponding aerodynamic coefficients. This model is able to predict $C_{l}$ and $C_{d}$ based on the input airfoil coordinates. Using NeuralFoil, we can quickly assess the performance characteristics of a large number of airfoil profiles without the need for time-consuming computational fluid dynamics (CFD) simulations [20].
@@ 3.4 Conditional Diffusion
While standard diffusion models are effective at generating novel airfoil samples, they typically lack mechanisms to ensure that these generated samples adhere to specific geometric or performance-based criteria. To address this limitation, we incorporate a conditional diffusion model that allows the generation process to be steered towards desired outcomes based on conditioning parameters, such as geometric features or aerodynamic performance metrics [29, 30, 6, 7].
@@ 3.4.1 Conditioning Mechanism
The core idea behind the conditional diffusion model is to integrate additional information, such as specific geometric constraints or aerodynamic coefficients, directly into the generative process. This conditioning is achieved by processing the conditioning data (e.g., desired lift coefficient $C_{l}$, drag coefficient $C_{d}$, maximum thickness, etc.) through a set of linear transformations, which map the conditioning information into a high-dimensional embedding space. Specifically, the conditioning information is passed through two fully connected (linear) layers followed by a ReLU activation function to produce a conditioning embedding.
@@
This embedding is then fused with the timestep information (representing the current state of the diffusion process) using a bilinear layer. The bilinear layer combines the time and conditioning embeddings into a single representation that captures both the progression of the diffusion process and the target constraints.
3.4.2 Integration into the U-Net Architecture
To effectively leverage the conditioning information during the generation process, we employ a Conditional U-Net architecture. The U-Net is a well-suited architecture for this task due to its ability to capture and retain spatial information at various scales through its encoder-decoder structure, enhanced with skip connections that prevent the loss of fine-grained details.
In our Conditional U-Net, the fused embedding (containing both the conditioning and timestep information) is injected into each block of the U-Net, ensuring that the model consistently takes the conditioning criteria into account throughout the diffusion process. This design enables the model to learn the relationship between the airfoil’s geometry and the specified conditioning parameters, allowing for the generation of airfoils that meet specific aerodynamic or geometric requirements.
As the diffusion model iteratively refines the airfoil from a noisy sample, the conditioning information guides the model to prioritize designs that align with the target specifications. This method not only improves the relevance of the generated airfoils but also enhances the model’s ability to explore feasible designs within the conditioned space.
3.4.3 Conditioning on Aerodynamic Coefficients
To further enhance the control over the generated airfoil profiles, we condition the model on aerodynamic coefficients such as $C_{l}$ and $C_{d}$, which are predicted using the NeuralFoil tool [20]. By incorporating these performance metrics directly into the conditioning mechanism, our model can generate airfoils tailored to specific aerodynamic requirements, such as maximizing lift or minimizing drag under certain operating conditions.
The inclusion of these aerodynamic coefficients as conditioning parameters adds another layer of flexibility to our diffusion model, enabling the generation of airfoils that not only adhere to geometric constraints but also meet specific performance targets. This capability is particularly valuable in practical design applications where achieving optimal aerodynamic performance is paramount.
3.4.4 Summary
The conditional diffusion model developed in this work provides a powerful and flexible framework for generating airfoil profiles that meet specific aerodynamic criteria. By integrating classifier-free guidance and conditioning on performance metrics, our approach addresses the limitations of traditional diffusion models, offering a robust solution for aerodynamic shape optimization.
3.5 Shape Similarity Measurement
To evaluate the similarity between the generated airfoils and the training dataset, we utilized the Chamfer distance [31, 32, 33], a metric commonly used in shape analysis to quantify the dissimilarity between two sets of points. The Chamfer distance is calculated by summing the average of the minimum Euclidean distances from each point in the first set to the nearest point in the second set, and vice versa. Mathematically, the Chamfer distance between two point sets $A$ and $B$ is defined as:
$d_{\text{Chamfer}}(A,B)=\frac{1}{|A|}\sum_{a\in A}\min_{b\in B}\|a-b\|_{2}+\frac{1}{|B|}\sum_{b\in B}\min_{a\in A}\|b-a\|_{2}$ (10)
A lower Chamfer distance indicates a higher similarity between the airfoils, while a higher distance suggests greater diversity. The distances were calculated using PyTorch on a GPU to efficiently handle the large quantity of computations involved.
@@ 4.2 Conditional Generation
Analysis of the relationship between the conditioned and generated values was conducted by plotting scatter plots of the input conditioning values against the NeuralFoil-evaluated results (Fig. 4 (b), (a)). The mean absolute error (MAE) was calculated to quantify the accuracy of the conditioning. The $C_{d}$-conditioned model achieved an MAE of 0.0018, whereas the $C_{l}$-conditioned model had a higher MAE of 0.04. The slope of the best-fit line for the $C_{d}$-conditioned model was 0.07, compared to 0.25 for the $C_{l}$-conditioned model, indicating that the $C_{d}$-conditioned model’s outputs were less sensitive to the input conditioning values. As the $C_{l}$ conditioning values exceeded the values seen in the UIUC samples, the $C_{l}$ conditioned model generations produced a greater range of $C_{l}$ evaluation values from generations, however the model still tended to produce airfoils with greater $C_{l}$ values corresponding to increasing conditioning values. This effect is likely due to passing conditioning values to the model that were not seen in training samples.
The range of output $C_{d}$ values was 0.0033 to 0.0226, and the range of output $C_{l}$ values was -0.3319 to 0.6068, compared to the input conditioning ranges of 0.038 to 0.0814 and -0.1701 to 3.21, respectively. The bias observed in the conditional diffusion model, particularly the narrower output range for $C_{l}$ values, may be attributed to the distribution of samples in the UIUC dataset, where most airfoils have $C_{l}$ values between -0.1 and 0.6, and $C_{d}$ values between 0.002 and 0.015. The model’s tendency to generate airfoils with more typical characteristics rather than extreme outliers suggests a limitation in exploring the full range of possible airfoil geometries.
@@ 4.2.1 Aerodynamic Performance
Figure 5(a) presents a scatter plot of the evaluated $C_{d}$ and $C_{l}$ values for both the $C_{d}$-conditioned and $C_{l}$-conditioned models. The $C_{d}$-conditioned model exhibits a broader range of $C_{d}$ values, demonstrating its ability to generate airfoils with diverse drag characteristics. However, it produces a narrower range of $C_{l}$ values, indicating more constrained lift performance. Conversely, the $C_{l}$-conditioned model generates a broader spectrum of $C_{l}$ values, showcasing its capacity to explore varied lift characteristics, while producing airfoils with a narrower range of $C_{d}$ values, favoring lower drag. This bias towards lower drag in the $C_{l}$-conditioned model is advantageous for maximizing the lift-to-drag ratio, making it particularly effective in producing efficient airfoil designs.
@@
4.2.2 Novelty of Generation
Figure 6: Boxplots of the Chamfer Distance between the UIUC dataset samples and the $C_{d}$ and $C_{l}$ conditioned models. Both models’ chamfer distance is predominantly around 0.005 to 0.025, with outliers as great as 0.08. The $C_{d}$ model exhibits slightly greater chamfer distances than the $C_{l}$ model.
To assess the novelty of the generated airfoils, we computed the Chamfer Distance between the profiles generated by the $C_{l}$- and $C_{d}$-conditioned models and their closest matching profiles from the UIUC airfoil dataset. As illustrated in Figure 6, the distribution of Chamfer Distances for both models highlights the extent of deviation from existing profiles in the dataset. This metric is crucial in capturing geometric dissimilarities between two sets of points, where a higher Chamfer Distance signifies a greater deviation or novelty, and a distance of zero would indicate an identical match.
To interpret the Chamfer Distance in the context of this study, we first examine the range of y-coordinates across both datasets. In the UIUC dataset, the minimum and maximum y-values range from -0.206 to 0.273, whereas for the $C_{d}$ and $C_{l}$-conditioned models, the corresponding ranges are -0.019 to 0.123 and -0.083 to 0.135, respectively. Given these ranges, we establish that a Chamfer Distance greater than 0.5 would indicate substantial novelty, as it exceeds the maximum point-to-point variation within any dataset.
For both models, Chamfer Distance values predominantly fall between 0.005 and 0.025, with outliers reaching up to 0.08. These values represent 1 to 16 percent of the maximal variation observed in the UIUC dataset, implying that even the most conservative samples exhibit at least 1 percent deviation from the training data. In contrast, the most novel generated airfoils demonstrate a geometric deviation of up to 16 percent, suggesting a meaningful degree of novelty while remaining within a physically plausible range.
The slightly elevated Chamfer Distances observed in the $C_{d}$-conditioned model relative to the $C_{l}$-conditioned model suggest that the $C_{d}$ model explores a broader geometric space. This broader exploration may stem from the wider range of drag values available in the training set, enabling the model to generate more varied and potentially unconventional shapes. Conversely, the $C_{l}$-conditioned model, while still producing novel airfoil profiles, appears more conservative, likely due to the narrower range of lift characteristics it is conditioned on.
Overall, these results confirm that the conditional diffusion models generate airfoil designs that deviate from the training dataset, introducing novel shapes. However, the generated airfoils remain within a viable aerodynamic design space, offering a balance between innovation and practicality. The higher novelty observed in the $C_{d}$ model suggests that conditioning on drag may facilitate greater exploration of the design space, making it a useful tool for discovering unique and potentially more efficient airfoil geometries.

### v3

@@ Abstract
The design of aerodynamic shapes, such as airfoils, has traditionally required significant computational resources and relied on predefined design parameters, which limit the potential for novel shape synthesis. In this work, we introduce a data-driven methodology for airfoil generation using a diffusion model. Trained on a dataset of preexisting airfoils, our model can generate an arbitrary number of new airfoils from random vectors, which can be conditioned on specific aerodynamic performance metrics such as lift ($C_{l}$) or drag ($C_{d}$). Our results demonstrate that the diffusion model effectively produces airfoil shapes with realistic aerodynamic properties, offering improvements in efficiency and flexibility. The model produces within the design space, and additionally generates new samples in sparse regions of the design space, with the potential to find new designs that exhibit favorable performance characteristics.
@@ 1 Introduction
The design and discovery of aerodynamic shapes, particularly airfoils, are fundamental challenges in aerospace engineering. Traditionally, the process of designing airfoils has relied heavily on computationally intensive simulations to explore the vast design space [1, 2]. This approach typically involves defining specific design parameters that can limit the ability to synthesize truly novel shapes [3]. As the demand for higher performance and more efficient aerodynamic structures increases, there is a growing need for advanced methodologies that can streamline the design process and facilitate the generation of innovative airfoil geometries [4].
Recent advances in machine learning have opened new avenues for the design and optimization of complex systems [5]. Data-driven approaches, such as generative models, have shown promise in various fields by learning from existing data to generate new high-quality samples[6, 7]. In the context of airfoil design, these models can potentially revolutionize the way we approach shape generation by learning the underlying distribution of airfoil geometries and their aerodynamic properties [3, 8, 9].
@@
The diffusion model, once trained, allows for the generation of new airfoil shapes from random vectors. This capability enables the synthesis of new airfoils that can be selected on the basis of desired Cl or Cd values. Our approach offers a data-driven framework that not only enhances the efficiency of the airfoil design process, but also has the potential to discover novel shapes with superior aerodynamic performance within the training dataset design space.
@@ 2.1 Geometry Parameterization
Traditional airfoil design has predominantly relied on computational simulations and fixed design parameters, which can restrict the exploration of innovative aerodynamic shapes [10]. These conventional methods, though effective in specific scenarios, often limit the ability to fully explore the expansive design space of aerodynamic structures [4]. This limitation arises primarily from the need for predefining the design space using specific shape parameters, such as camber and thickness, which can constrain the range of possible geometries that can be considered during the airfoil design process.
To address these limitations, there has been a shift towards advanced machine learning techniques that allow for more flexible and efficient exploration of the design space. These approaches, including techniques like Bézier and B-spline parameterizations, offer greater control over airfoil geometry by representing shapes with a set of control points that can be adjusted to generate airfoil coordinate and continuous surfaces. Such methods have been integrated into optimization frameworks to enhance the precision of geometric design while maintaining smoothness and manufacturability [11, 12]. However, while these traditional geometric parameterizations provide a level of flexibility, they still require significant manual intervention and are often limited in handling highly complex shapes, especially in high-dimensional design spaces.
@@
Moreover, Diffusion Models, as explored in recent studies, represent a significant advancement in airfoil geometry parameterization. These models operate within a learned latent space, allowing for the generation of airfoil shapes with fewer training samples and maintaining high-quality outputs [14]. The ability of Diffusion Models to integrate specific design constraints directly into the generative process provides a powerful tool for generating aerodynamic shapes in a data-efficient manner.
As the field continues to evolve, hybrid approaches that combine traditional geometric parameterization with advanced machine learning techniques are emerging. These methods aim to leverage the strengths of both worlds—preserving the airfoil coordinateness and manufacturability of traditional parameterizations while exploiting the flexibility and data-driven insights of deep learning models. The ongoing research in this area focuses on refining these techniques to accommodate increasingly complex aerodynamic and geometric constraints, ultimately pushing the boundaries of what is possible in airfoil design.
@@ 2.2 Deep Learning Approaches
Generative models, including Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs), have shown considerable promise in advancing airfoil design by overcoming the limitations of traditional design methods. For instance, the Airfoil GAN model [3] encodes existing airfoil shapes into latent vectors, enabling the generation of novel airfoils with specific aerodynamic properties, without relying on predefined design parameters. This approach not only facilitates the exploration of high-dimensional design spaces but also integrates a genetic algorithm for optimizing aerodynamic performance, allowing the synthesized airfoils to evolve towards desired properties. Moreover, the model’s ability to generate airfoil coordinate and realistic airfoils directly from the latent space, coupled with its interpretability through feature clustering, sets it apart from traditional techniques. In comparison, the use of free-form deformation generalized adversarial networks (FFD-GAN) [8] has advanced the parameterization of three-dimensional aerodynamic shapes, yet it often requires predefined design formulations, limiting its flexibility in exploring novel geometries. The Airfoil GAN’s potential for application to other aerodynamic and engineering domains further underscores its significance in the field of design optimization.
Despite their potential, GAN-based methods often face challenges related to training stability and the need for large datasets. Addressing these issues, DiffAirfoil [15] introduces a novel approach that leverages a diffusion model within a learned latent space, specifically tailored for aerodynamic shape optimization. This model significantly enhances data efficiency, requiring substantially fewer training geometries compared to GANs, while still generating high-quality and diverse airfoil samples. The diffusion process not only ensures stable training and robust sampling, even under data-scarce conditions, but also mitigates the common pitfalls associated with adversarial training, such as mode collapse [16, 17].
@@ 3.2.2 NeuralFoil Overview
NeuralFoil leverages a deep learning model trained on a comprehensive dataset of airfoil shapes and their corresponding aerodynamic coefficients. This model is able to predict $C_{l}$ and $C_{d}$ based on the input airfoil coordinates. While some skepticism may arise regarding whether a neural network based model can adhere to physical principles, NeuralFoil addresses this concern by using a physics-informed neural network (PINN) architecture trained on tens of millions of XFOIL simulations [22]. This ensures that the model respects aerodynamic laws while providing rapid predictions. Furthermore, NeuralFoil has been benchmarked against XFOIL, achieving accuracy within a few percent of XFOIL’s predictions while offering significant computational speedups [22]. Additionally, NeuralFoil provides an “analysis confidence” output for each prediction, quantifying uncertainty and ensuring robust design optimization [22].
@@ 3.4 Conditional Diffusion
Standard diffusion models are effective at generating novel airfoil samples but often lack mechanisms to ensure that these samples meet specific geometric or performance-based criteria. To address this, we incorporate a conditional diffusion model that steers the generation process toward desired outcomes based on conditioning parameters, such as geometric features or aerodynamic performance metrics [31, 32, 6, 7].
@@ 3.4.1 Conditioning Mechanism
The conditional diffusion model integrates additional information, such as geometric constraints or aerodynamic coefficients, directly into the generative process. The conditioning data (e.g., lift coefficient $C_{l}$, maximum thickness, or maximum camber) is processed through linear transformations to map it into a high-dimensional embedding space. This is followed by a ReLU activation function.
@@
The resulting conditioning embedding is fused with timestep information using a bilinear layer, combining both into a single representation that captures the progression of the diffusion process and target constraints.
3.4.2 Conditional U-Net Architecture for 1D Data
We employ a Conditional U-Net architecture to incorporate conditioning information during airfoil generation. The U-Net’s encoder-decoder structure, enhanced with skip connections, is well-suited for retaining spatial information at various scales. In our case, the U-Net is adapted for 1D, two-channel input data (e.g., airfoil coordinates), replacing the traditional 2D convolutions with 1D convolutions for downsampling and upsampling operations.
The input data is first processed by an initial 1D convolutional layer to match the network’s dimensionality requirements. The downsampling path consists of multiple stages, each containing ResNet blocks, linear attention layers, and downsampling operations (either strided or learned convolutions). These layers progressively reduce the spatial dimensions while increasing feature channels. The upsampling path mirrors this structure but progressively increases spatial dimensions while reducing feature channels. The final output is produced by residual blocks and a 1D convolutional layer.
To effectively integrate conditioning information (such as desired aerodynamic coefficients $C_{l}$ and $C_{d}$), we fuse the conditioning data with timestep embeddings at each stage of the U-Net. Time step information is encoded via sinusoidal embeddings processed through a multi-layer perceptron (MLP), while conditional inputs are processed through linear layers and bilinear layers to generate conditional embeddings. These embeddings are combined with time embeddings and injected into ResNet blocks throughout both downsampling and upsampling paths.
Mathematically, input data is encoded into a latent representation via the U-Net’s downsampling path. The combined time and conditional embeddings are used to condition ResNet blocks across the network, ensuring that generated outputs satisfy specified conditions. The latent representation is then decoded back into structured data through upsampling layers, guiding the model toward airfoil designs that meet target specifications.
@@ 4 Results and Discussion
4.1 Generated Profile Smoothing
To address the lack of guaranteed smoothness in airfoils generated by the diffusion model, we implement a smoothing framework based on B-spline interpolation, which offers a smooth approximation of the airfoil shape [33, 34]. This is achieved using the BSpline and splrep functions provided by the SciPy Python library [35].
To ensure compatibility with this method, we preprocess the airfoil coordinates by separating the upper and lower surfaces and reordering the points to have increasing $x$ values. For the upper surface, where $x$ values are typically decreasing, we reorder the points as:
$\{(x_{i},y_{i})\}_{i=1}^{n}\rightarrow\{(x_{\sigma(i)},y_{\sigma(i)})\}_{i=1}^{n},$
where $\sigma$ is a permutation such that $x_{\sigma(1)}<x_{\sigma(2)}<\dots<x_{\sigma(n)}$.
After smoothing, the B-spline representation is restored to the original orientation to preserve the original data format while achieving a smoother representation.
4.2 Shape Similarity Measurement
To evaluate the similarity between the generated airfoils and the training dataset, we utilized Principal Component Analysis (PCA), a dimensionality reduction technique that transforms high-dimensional data into a lower-dimensional space while preserving the most significant modes of variation [36, 37]. PCA computes linear combinations of the original variables, called principal components, which are uncorrelated and ranked by the variance they explain in the data.
The PCA process begins by centering the observation matrix $\mathbf{X}$, where each row represents an airfoil, by subtracting the mean of each feature. The variance-covariance matrix $\mathbf{S}$ is then constructed, and the principal components $\mathbf{V}$ are calculated as:
$\mathbf{V}=\mathbf{A}^{\top}\mathbf{X},$
where $\mathbf{A}$ is a matrix of eigenvectors of $\mathbf{S}$, and the eigenvalues $\lambda_{i}$ represent the variance captured by the corresponding principal components. For this study, PCA was applied to the training dataset to project airfoil shapes into a reduced-dimensional space. The same transformation was then applied to the generated airfoils, enabling direct comparisons in this space.
@@ 4.4 Conditional Generation
Analysis of the relationship between the conditioned and generated values was conducted by plotting scatter plots of the input conditioning values against the NeuralFoil-evaluated results (Fig. 4 (b), (a)). The mean absolute error (MAE) was calculated to quantify the accuracy of the conditioning. The $C_{d}$-conditioned model achieved an MAE of 0.0018, whereas the $C_{l}$-conditioned model had a higher MAE of 0.04. The slope of the best-fit line for the $C_{d}$-conditioned model was 0.07, compared to 0.25 for the $C_{l}$-conditioned model, indicating that the $C_{d}$-conditioned model’s outputs were less sensitive to the input conditioning values. As the $C_{l}$ conditioning values exceeded the values seen in the UIUC samples, the $C_{l}$ conditioned model generations exhibited a greater standard deviation, however the model still tended to produce airfoils with greater $C_{l}$ values corresponding to increasing conditioning values. This effect is likely due to passing conditioning values to the model that were not seen in training samples.
The range of output $C_{d}$ values was 0.0033 to 0.0226, and the range of output $C_{l}$ values was -0.3319 to 1.09. These outputs were narrower compared to the input conditioning ranges of 0.038 to 0.0814 for $C_{d}$ and -0.1701 to 3.21 for $C_{l}$. The maximal $C_{l}$ value observed in the UIUC dataset is approximately 1.25; however, this is an outlier. The majority of airfoils in the dataset exhibit $C_{l}$ values between 0.0 and 0.75, suggesting that the model is sufficiently capable of representing most samples in the dataset, excluding the most extreme cases.
To approximate a one-to-one correspondence between input conditioning values and output airfoil performance coefficients, the input values can be scaled by a constant factor equal to the inverse slope of the line of best fit. This adjustment ensures alignment, provided the desired conditioning values remain within the bounds supported by the model.
@@ 4.4.1 Aerodynamic Performance
We evaluated the performance of generated airfoils using the NeuralFoil library [22], with angle of attack set to 0 degrees, Reynolds number of $1\times 10^{6}$ and mach number 0.0. Figure 5(a) presents a scatter plot of the evaluated $C_{d}$ and $C_{l}$ values for both the $C_{d}$-conditioned and $C_{l}$-conditioned models. The $C_{d}$-conditioned model exhibits a broader range of $C_{d}$ values, demonstrating its ability to generate airfoils with diverse drag characteristics. However, it produces a narrower range of $C_{l}$ values, indicating more constrained lift performance. Conversely, the $C_{l}$-conditioned model generates a broader spectrum of $C_{l}$ values, showcasing its capacity to explore varied lift characteristics, while producing airfoils with a narrower range of $C_{d}$ values, favoring lower drag. This bias towards lower drag in the $C_{l}$-conditioned model is advantageous for maximizing the lift-to-drag ratio, making it particularly effective in producing efficient airfoil designs.
@@
4.4.2 Design Space Analysis
To assess the overlap and exploration of the design space, we applied PCA to both the training dataset and the generated airfoils conditioned on $C_{l}$ and $C_{d}$. Figure 6 shows the PCA projections of the UIUC dataset alongside the generated samples.
Figure 6: PCA projections of the UIUC dataset samples compared with the principal components of the $C_{l}$-conditioned model (Figure a) and the $C_{d}$-conditioned model (Figure b). The principal components of both models align closely with the most frequently occurring projections in the UIUC dataset. The $C_{l}$-conditioned model projections encompass the UIUC projections uniformly, while the $C_{d}$-conditioned model projections also encompass the dataset but are noticeably shifted to the left relative to the center of the UIUC distribution.
From Figure 6, the $C_{l}$-conditioned model (Figure 6(a)) is centered around the most frequently occurring projections of the UIUC dataset, with a slight expansion in the space covered. In contrast, the $C_{d}$-conditioned model (Figure 6(b)) demonstrates a shift to the left, indicating an exploration of different characteristics of the design space compared to the $C_{l}$-conditioned model.
This result suggests that the airfoil diffusion model adheres to the design space defined by the training dataset, ensuring physically feasible airfoil designs. Simultaneously, the model exhibits the capability to generate novel geometries within this realistic design space, offering potential for discovering new airfoil profiles with favorable performance characteristics. This feature underscores the utility of the diffusion model in assisting airfoil design exploration and innovation.